Suppose an economy's real GDP is $14,736 in year 1 and $39,693 in year 2. Also suppose that the same economy's population is 174 in year 1 and 271 in year 2

Calculate the growth rate of the economy's real GDP?

Calculate the growth rate of the economy's real GDP per capita
1
Expert's answer
2012-10-17T09:15:33-0400
GDP growth rate = $39,693/$14,736 = 2.694 = 169.4%,
GDP per capita growth
rate = ($39,693/271)/($14,736/174) = 146.47/84.69 = 1.729 = 72.9%,